Find the percent composition of each element in caffeine, C8H10N4O2
1
Expert's answer
2018-04-07T08:15:13-0400
At first find the molar mass of each element in caffeine:
M(C) = 12 g/mol x 8 = 96
M(H) = 1 g/mol x 10 = 10
M(N) = 14 g/mol x 4 = 56
M(O) = 16 g/mol x 2 = 32
M(C8H10N4O2) = 96 + 10 + 56 + 32 = 194 g/mol
Now find the percent composition of each element:
w = (molar mass of component/molar mass of total) x 100%
w(C) = (96/194) x 100 = 49.48%
w(H) = (10/194) x 100 = 5.15%
w(N) = (56/194) x 100 = 28.87%
w(O) = (32/194) x 100 = 16.50%